What is an equation of the line that passes through the point (−1,−3) and is parallel to the line 6x−y=4. Express your answer in slope-intercept form.
Answer by mathstutor458(57)

line 6x−y=4.
y=6x-4
slope m=6 (compare with y=mx+b)
point is (-1,-3)
Use y = mx+b and the point to find b.
-3=6*(-1)+b
-3=-6+b
b=6-3=3
Therefore, y=6x+3 ---->answer